Solve for t.
3000(1-e^(-2t))=1,200
So,I know you divide everything by 3000,right?
1-e^(-2t)=2/5
Now what do I do?

Answer by Alan3354(69443)

Solve for t.
3000(1-e^(-2t))=1,200
So,I know you divide everything by 3000,right?
1-e^(-2t)=2/5
Now what do I do?
=====================
You solve for t.
1-e^(-2t) = 0.4
-e^(-2t) = 0.4 - 1 = -0.6
e^(-2t) = 0.6
ln(e^(-2t)) = ln(0.6)
-2t = ln(0.6)
t = -ln(0.6)/2
t =~ 0.2554